# Optical turbulence profiling at the Table Mountain Facility with the Laser Communication Relay Demonstration GEO downlink

### v1

@@ 3.1 Dual $r_{0}$ comparison
Figure 5 highlights the relation between $r_{0,\text{scint}}$ and $r_{0,\text{sector}}$ over a range of turbulence strengths with measurements conducted in daytime and nighttime at TMF. Although Figure 5 displays a robust relation between these measurement techniques, there is a significant trend of $r_{0,\text{sector}}$ overestimating turbulence with respect to $r_{0,\text{scint}}$. $r_{0,\text{sector}}$ is typically expected to be larger than $r_{0,\text{scint}}$, as described in [1], and the opposite result here could be caused by numerous issues such as strong upper layer turbulence, astigmatism, or error in the expected pixel scale. Pixel scale for this system cannot be easily measured due to the narrow linewidth filter blocking all stars, i.e. plate-solving or binaries cannot be observed to measure pixel scale. If the system is not at the ideal focal length, this could also cause the larger than expected ring widths we observe. Further observations are required to constrain these relations and propose explanations. No correlation was observed between upper atmosphere $C_{n}^{2}$ and this ratio over the preliminary dataset, so we choose not to reproduce that plot here. Another possible cause of this is how each method differs in response to deviations from the weak-scintillation regime, so more data in weaker turbulence could be used to assess this theory [1, 14]. Nonetheless we consider the strong correlation between $r_{0}$ methods a good first indicator the instrument is producing reliable results.
@@ 4 Accuracy analysis
We compare against three turbulence monitors at TMF to further assess RINGSS. Comparing turbulence monitors is often difficult due to different measurement techniques, positions on the mountain, and from the stochastic nature of turbulence, nonetheless this remains the best tool available [20]. TMF hosts a solar scintillometer, Polaris monitor, and boundary-layer scintillometer, the latter being covered separately in section 5 [21, 22]. The solar scintillometer was previously calibrated against a solar DIMM and is mounted alongside the Polaris monitor on the roof of OCTL, approximately 50 m from RINGSS (see Figure 1) [22]. All instruments, including RINGSS, are airmass-corrected to represent values at zenith for appropriate comparison.
A nearly continuous observing block from the 24th of May 2023 is plotted in Figure 6 to highlight the comparison with the solar scintillometer and Polaris monitor, in addition to showing the diurnal evolution of turbulence. Residuals are computed by time-averaging all methods into two-minute bins. Gaps of RINGSS data in Figure 6 are due to technical issues and pauses in LCRD operations. There is also a break between when the solar scintillometer operates and the Polaris monitor is able to begin measurements. Figure 6 shows that two different $r_{0}$ estimates from RINGSS agree with two separate instruments over the presented time range, and that the technique is reliable in strong daytime turbulence. Provided the scintillation spectrum saturates in strong turbulence, presenting a major challenge for profilers which rely on weak-scintillation assumptions, the performance we see in daytime conditions is encouraging. Turbulence is significantly weaker at $1550$ nm than typical observations in the visible, which is undoubtedly a major factor for RINGSS measurements remaining reliable, in addition to the saturation-correction we apply as described in [1] and developed in [14].
@@
We then compare between all instruments over the entire dataset, time-averaging to 10 minute bins for comparison as shown in Figure 7. Although any degree of time-averaging is necessary for comparison, 10 minute bins are chosen partly because these instruments are not in the same position on the mountain, and not observing through the same channel, hence we don’t expect short term agreement ($\lesssim 1$ minute). Further, we choose 10 minutes as an intuitive yet arbitrary scale to characterise current conditions by given the stochastic nature of turbulence. Figure 7 shows the Pearson correlation coefficients ($R$) and linear regression fits between RINGSS, the Polaris monitor, and solar scintillometer over this preliminary dataset. We observe reasonable correlation coefficients, i.e $R>0.7$ for all cases, promising in the context of turbulence monitors, but large systematic offsets ($\beta$) and slopes all significantly less than $1$. Nighttime RINGSS observations tends to agree more with the Polaris monitor than daytime RINGSS with the solar scintillometer. This could be either due to RINGSS being more accurate in weaker turbulence, e.g. due to insensitivity or increased error towards a strong boundary layer, or that the Polaris monitor is more accurate than the solar scintillometer. In all cases we observe RINGSS estimating stronger turbulence than alternate instruments, though more data is required to assess whether this is systematic. Long term operation of RINGSS at TMF is required to see if these early promising correlations improve and hold over a broader range of conditions. Inherently different measurement techniques, instrument error, and different positions on TMF are all reasons to not expect a robust agreement between instruments in the long-term that might be expected of other meteorological parameters.
@@ 5 Boundary layer measurements
Turbulence will typically be worst in the daytime, where a convective boundary layer can form and dominate turbulence, so sensitivity to this layer is critical [23]. We are therefore interested in determining the accuracy of RINGSS near the surface in daytime, i.e. limiting turbulence conditions in daytime operation. To assess RINGSS accuracy in this regime we compare with a boundary layer scintillometer (BLS) which is installed at OCTL to sample $C_{n}^{2}$ at high frequency in a horizontal channel at all times of day [21]. The BLS measures turbulence over a valley to a nearby mountain from TMF in the same direction that RINGSS observes LCRD, though its location at OCTL is raised off the ground unlike RINGSS. We compare BLS measurements with RINGSS turbulence boundary layer estimates by dividing the lowest $J_{i}$ integral by layer height, i.e. from 00 m to $250$ m, to estimate $C_{n}^{2}$. This method assumes a uniform distribution of $C_{n}^{2}$ through the layer but given RINGSS is observing over a steep valley and at a slant path we cannot justify a more sophisticated solution.
Figure 8 shows the distribution of $C_{n}^{2}$ from the BLS compared with the ground layer $C_{n}^{2}$ estimate from RINGSS. We observe agreement within expectations for differing instruments measuring $C_{n}^{2}$, though with heteroscedasticity towards small $C_{n}^{2}$ likely due to a lack of data and different paths through the boundary layer. Nonetheless, Figure 8 demonstrates that RINGSS is capable of approximately estimating the boundary layer turbulence even in daytime. However, it should be noted that this preliminary dataset represents fairly weak boundary layer turbulence, even at daytime, likely due to the steep valley, altitude of TMF, and cool weather. Significantly more data is required over a substantial dynamic range to established a robust correlation with RINGSS and the BLS, particularly for $C_{n}^{2}>10^{-15}$. Correlation of boundary layer turbulence with wind-speed and direction from meteorological instruments will also be considered in future work.
@@ 6 Profile results
Turbulence profiling also allows us to estimate a number of important integrated parameters, particularly important for AO operation at OCTL during LCRD links. Table 1 shows the average atmospheric coherence time ($\tau_{0}$), Greenwood frequency ($f_{G}$) and isoplanatic angle ($\theta_{0}$) from all measured profiles in the preliminary dataset, split into daytime and nighttime. Estimating $\tau_{0}$ and $f_{G}$ requires knowledge of the wind speed profile, though the turbulence-weighted wind velocity, $\bar{V}$, can be substituted for some loss of accuracy [25]. Equation 5 shows how $\tau_{0}$ and $f_{G}$ are computed for $\bar{V}$ if wind speed profiling, e.g. measured by a Stereo-SCIDAR, is not available [26].
@@
Figure 9: Turbulence profiles measured by RINGSS at the Table Mountain Facility, California. Day and night time profiles are split to highlight the diurnal evolution of turbulence, particularly in the boundary layer. $J_{i}$ values refer to discrete layers of turbulence representing integrals of $C_{n}^{2}dh$.
To estimate $\bar{V}$ with RINGSS we instead use the atmospheric second moment of the wind, $V_{2}$, described by Kornilov in [27] with implementation for RINGSS described in [1]. $V_{2}$ is principally measured with RINGSS from the defocus velocity, a concept with origins in the FADE instrument, and has been shown to follow $V_{5/3}$ closely, which defines $\tau_{0}$ [15]. We do not reproduce this methodology here, as it is involved and best explained by Kornilov [25, 15, 27]. Measuring $V_{2}$ and the defocus velocity is sensitive to blurring so the short exposure time, $200$ µs, afforded by the strong LCRD downlink is significant in improving accuracy here.
Table 1: Integrated parameters from RINGSS turbulence profile measurements at TMF. Data from short preliminary dataset is split into daytime and nighttime LCRD links. Uncertainties are the standard deviation from all parameter calculations over the dataset
$r_{0,\text{scint}}$ [cm] $r_{0,\text{sector}}$ [cm] $f_{G}$ [Hz] $\theta_{0}$ [arcsec] $\tau_{0}$ [ms]
Night $7.2\pm 1.8$ $6.0\pm 1.6$ $55\pm 19$ $1.7\pm 0.5$" $2.7\pm 0.8$
Day $5.0\pm 1.4$ $4.3\pm 1.2$ $68\pm 33$ $2.1\pm 0.4$ $2.2\pm 0.4$
In Table 1 we see almost all parameters degrade in daytime compared with nighttime, however $\theta_{0}$ worsens slightly at nighttime due to the increased turbulence above 2 km visible in Figure 9. This highlights one of the benefits of profiling over integrated turbulence measurements, i.e. $\theta_{0}$ is an important parameter to consider for uplink pre-compensation with AO which is driven by high altitude turbulence [28].
@@ 7 Conclusion
We have successfully demonstrated the deployment of a RINGSS turbulence profiler, an instrument initially demonstrated in 2021, and conducted the first turbulence profiling with a laser communications satellite using the LCRD GEO optical downlink signal [1]. This is also the first time that a RINGSS instrument has been used in daytime and strong turbulence conditions, highlighting the measurement technique is robust over a broad range of conditions. Preliminary results have been compared with a suite of other turbulence monitors at TMF, showing the reasonable correlations despite the small dataset and difficulties of finding agreement among turbulence instruments. Efforts are also ongoing to automate this instrument with closed-loop tracking, auto-focusing, and an observatory control system. While we currently lack the data to make broad and robust statements on the instrument performance and site characteristics we have demonstrated the unique concept of FSOC turbulence profiling which could improve the performance of GEO feeder networks. Future operations of RINGSS at TMF alongside LCRD operations will aim to constrain the correlations we have presented here and develop this instrument into a capability for improving reliability with short-term turbulence prediction and inform the AO control system on OCTL.

### v2

@@ 3.1 Dual $r_{0}$ comparison
Figure 5 highlights the relation between $r_{0,\text{scint}}$ and $r_{0,\text{sector}}$ over a range of turbulence strengths with measurements conducted in daytime and nighttime at TMF. Although Figure 5 displays a robust relation between these measurement techniques, there is a significant trend of $r_{0,\text{sector}}<r_{0,\text{scint}}$, i.e. overestimating turbulence strength. $r_{0,\text{sector}}$ is typically expected to be larger than $r_{0,\text{scint}}$, as described in [1], and the opposite result here could be caused by numerous issues we discuss here. Radial sector motion, like with a DIMM, can overestimate free atmosphere turbulence, an effect enhanced by astigmatism that is described in [15]. However, error in the expected pixel scale is more likely as it cannot be easily measured in this system due to the narrow linewidth filter blocking all stars, i.e. plate-solving or binaries cannot be observed to measure pixel scale. The larger than expected $R_{\text{width}}$ we observe also indicates an out of specification focal length and therefore pixel scale.
Another possible cause is the profile restoration underestimating turbulence due to scintillation saturation, so more data in weaker turbulence could be used to assess this theory [1, 15]. Nonetheless, further observations are required to draw conclusions, though we consider the strong correlation ($R=0.95$) between methods a good first indicator the instrument is producing reliable results.
@@ 4 Accuracy analysis
We compare against three turbulence monitors at TMF to further assess RINGSS. Comparing turbulence monitors is often difficult due to different measurement techniques, positions on the mountain, and from the stochastic nature of turbulence, nonetheless this remains the best tool available [21]. TMF hosts a solar scintillometer, Polaris motion monitor, and boundary-layer scintillometer, the latter being covered separately in section 5 [22, 23, 24]. The Polaris monitor is a single aperture instrument which derives $r_{0}$ from absolute image motion of Polaris captured at 50 Hz [24]. The solar scintillometer was previously calibrated against a solar DIMM and is mounted alongside the Polaris monitor on the roof of OCTL, approximately 50 m from RINGSS (see Figure 1) [23]. All instruments, including RINGSS, are corrected to represent values at zenith and 500 nm for appropriate comparison. These integrated seeing monitors, located on the roof of OCTL, are shown in Figure 6.
Figure 6: Integrated turbulence monitors at Table Mountain Facility. Left: Polaris motion monitor [24]. Right: solar scintillometer [23]. Note that the Polaris motion monitor points North, while RINGSS and solar scintillometer point South.
A nearly continuous observing block from the 24th of May 2023 is plotted in Figure 7 to highlight the comparison with the solar scintillometer and Polaris monitor, in addition to showing the diurnal evolution of turbulence. Residuals are computed by time-averaging all methods into two-minute bins. Gaps of RINGSS data in Figure 7 are due to technical issues and pauses in LCRD operations. There is also a break between when the solar scintillometer operates and the Polaris monitor is able to begin measurements. Figure 7 shows that two different $r_{0}$ estimates from RINGSS agree with two separate instruments over the presented time range, including in strong turbulence. The daytime performance we see is encouraging as scintillation saturation is a major challenge for profilers which rely on weak-scintillation assumptions. However, scintillation is significantly weaker at $1550$ nm than typical observations in the visible, which is a major factor for this instruments reliability (see section 6). We nonetheless still apply a saturation correction as described in [1] and developed in [15], though it’s impact on $r_{0}$ estimation is $<5\%$.
@@
We then compare between all instruments over the entire dataset, time-averaging to 10 minute bins for comparison as shown in Figure 8. Although any degree of time-averaging is necessary for comparison, 10 minute bins are chosen partly because these instruments are not in the same position on the mountain, and not observing through the same path, hence we don’t expect short term agreement ($\lesssim 1$ minute). Further, we choose 10 minutes as an intuitive yet arbitrary scale to characterise current conditions by given the stochastic nature of turbulence. Figure 8 shows the Pearson correlation coefficients ($R$) and linear regression fits between RINGSS, the Polaris monitor, and solar scintillometer over this preliminary dataset. We observe reasonable correlation coefficients, i.e $R>0.7$ for all cases, promising in the context of turbulence monitors, but large systematic offsets ($\beta$) and slopes all significantly less than $1$. Nighttime RINGSS observations tends to agree more with the Polaris monitor than daytime RINGSS with the solar scintillometer. This could be either due to RINGSS being more accurate in weaker turbulence, e.g. due to insensitivity or increased error towards a strong boundary layer, or that the Polaris monitor is more accurate than the solar scintillometer. In all cases we observe RINGSS estimating stronger turbulence than alternate instruments, though more data is required to assess whether this is systematic. Long term operation of RINGSS at TMF is required to see if these early promising correlations improve and hold over a broader range of conditions. Inherently different measurement techniques, instrument error, and different positions on TMF are all reasons to not expect a robust agreement between instruments in the long-term that might be expected of other meteorological parameters.
@@ 5 Boundary layer measurements
Effects of turbulence are typically worst in the daytime, where a convective boundary layer can form and dominate turbulence, so sensitivity to this layer is critical [25]. We are therefore interested in determining the accuracy of RINGSS near the surface in daytime, i.e. limiting turbulence conditions in daytime operation. To assess RINGSS accuracy in this regime we compare with a SCINTEC BLS900 boundary layer scintillometer (BLS), which is installed on the roof of OCTL (nearby to instruments shown in Figure 6) to sample $C_{n}^{2}$ at high frequency in a horizontal path at all times of day [22]. The BLS points south and measures turbulence over a valley to a nearby mountain $3.4$ km from TMF in the same direction that RINGSS observes LCRD, though its location at OCTL is raised off the ground unlike RINGSS. We compare BLS measurements with RINGSS turbulence boundary layer estimates by dividing the lowest $J_{i}$ integral by layer height, i.e. from 00 m to $250$ m, to estimate $C_{n}^{2}$. This method assumes a uniform distribution of $C_{n}^{2}$ through the layer but given RINGSS is observing over a steep valley and at a slant path we cannot justify a more sophisticated solution.
Figure 9 shows the distribution of $C_{n}^{2}$ from the BLS compared with the ground layer $C_{n}^{2}$ estimate from RINGSS. We observe agreement within expectations for differing instruments measuring $C_{n}^{2}$, though with heteroscedasticity towards small $C_{n}^{2}$ likely due to a lack of data and different paths through the boundary layer. Nonetheless, Figure 9 demonstrates that RINGSS is capable of approximately estimating the boundary layer turbulence even in daytime. However, it should be noted that this preliminary dataset represents fairly weak boundary layer turbulence, even at daytime, likely due to the steep valley, altitude of TMF, and cool weather. Significantly more data is required over a substantial dynamic range to established a robust correlation with RINGSS and the BLS, particularly for $C_{n}^{2}>10^{-15}$ m-2/3. Correlation of boundary layer turbulence with wind-speed and direction from meteorological instruments will also be considered in future work.
@@ 6 Profile results
Turbulence profiling also allows us to estimate a number of important integrated parameters, particularly important for AO operation at OCTL during LCRD links. Table 1 shows the median atmospheric coherence time ($\tau_{0}$), Greenwood frequency ($f_{G}$) and isoplanatic angle ($\theta_{0}$) from all measured profiles in the preliminary dataset, split into daytime and nighttime. Estimating $\tau_{0}$ and $f_{G}$ requires knowledge of the wind speed profile, though the turbulence-weighted wind velocity, $\bar{V}$, can be substituted for some loss of accuracy [27]. Equation 5 shows how $\tau_{0}$ and $f_{G}$ are computed for $\bar{V}$ if wind speed profiling, e.g. measured by a Stereo-SCIDAR, is not available [28]. We also include an estimation of the Rytov variance, $s_{0}^{2}$, as formulated in [29] for the small-signal approximation (RINGSS cannot directly measure the scintillation index like the solar scintillometer). $s_{0}^{2}$ is estimated for the operating wavelength of $1550$ nm and highlights the benefit of observing in infrared for remaining in the weak to moderate scintillation regime, i.e. $\sigma_{0}^{2}\approx 0.06$ as in Table 1 would be $\approx 0.2$ if observed at $550$ nm. It was shown in [1] that depending on the profile structure, even a moderate $s_{0}^{2}$ can be problematic for resolving layers close to the surface.
@@
Figure 10: Median turbulence profiles measured by RINGSS at the Table Mountain Facility, California. Day and night time profiles are split to highlight the diurnal evolution of turbulence, particularly in the boundary layer. Error bars are calculated from the median absolute deviation.
To estimate $\bar{V}$ with RINGSS we instead use the atmospheric second moment of the wind, $V_{2}$, described by Kornilov in [30] with implementation for RINGSS described in [1]. $V_{2}$ is principally measured with RINGSS from relating fast focus variation to the temporal structure function for atmospheric piston-modes, a concept with origins in the FADE instrument, and has been shown to follow $V_{5/3}$ closely, which defines $\tau_{0}$ [16]. We do not reproduce this methodology here, as it is involved and best explained by Kornilov [27, 16, 30]. Measuring $V_{2}$ from rapid changes in defocus caused by scintillation as $R_{\text{radius}}$ is sensitive to blurring so the short exposure time, $200$ µs, afforded by the strong LCRD downlink, is significant in improving accuracy here.
Table 1: Median integrated parameters from RINGSS turbulence profile measurements at TMF. Data from short preliminary dataset is split into daytime and nighttime LCRD links. Uncertainties are the standard deviation from all parameter calculations over the dataset
$r_{0,\text{scint}}$ [cm] $r_{0,\text{sector}}$ [cm] $f_{G}$ [Hz] $\theta_{0}$ [arcsec] $\tau_{0}$ [ms] $\sigma_{0}^{2}$$\>{}^{\dagger}$
Night $7.2\pm 1.8$ $6.0\pm 1.6$ $55\pm 19$ $1.7\pm 0.5$" $2.7\pm 0.8$ $0.063\pm 0.02$
Day $5.0\pm 1.4$ $4.3\pm 1.2$ $68\pm 33$ $2.1\pm 0.4$ $2.2\pm 0.4$ $0.064\pm 0.03$
  • Rytov variance, $\sigma_{0}^{2}$, is computed for the measurement wavelength of $1550$ nm, rather than $500$ nm like other integrated parameters [29].
In Table 1 we see almost all parameters degrade in daytime compared with nighttime, however $\theta_{0}$ worsens slightly at nighttime due to the increased turbulence above 2 km visible in Figure 10. This highlights one of the benefits of profiling over integrated turbulence measurements, i.e. $\theta_{0}$ is an important parameter to consider for uplink pre-compensation with AO which is driven by high altitude turbulence [31]. The Rytov variance approximation from [29] is also less sensitive to increases of turbulence strength near the surface which explains the very similar $\sigma_{0}^{2}$ values between day and night.
@@ 7 Conclusion
We have successfully demonstrated the deployment of a RINGSS turbulence profiler, an instrument initially demonstrated in 2021, and conducted the first turbulence profiling with a laser communications satellite using the LCRD GEO optical downlink signal [1]. This is also the first time that a RINGSS instrument has been used in daytime and strong turbulence conditions, highlighting the measurement technique is robust over a broad range of conditions. Preliminary results have been compared with a suite of other turbulence monitors at TMF, showing reasonable correlations ($R=0.75-0.86$) despite the small dataset and difficulties of finding agreement among turbulence instruments. Efforts are also ongoing to automate this instrument with closed-loop tracking, auto-focusing, and an observatory control system. While we currently lack the data to make broad and robust statements on the instrument performance and site characteristics we have demonstrated the unique concept of FSOC turbulence profiling which could improve the performance of GEO feeder networks. Future operations of RINGSS at TMF alongside LCRD operations will aim to constrain the correlations we have presented here and develop this instrument into a capability for improving reliability with short-term turbulence prediction and inform the AO control system on OCTL.
@@
Portions of this work were presented at the SPIE Photonics West Free-Space Laser Communications XXXVI conference in 2024 as “Measuring the vertical profile of atmospheric turbulence with the laser communication relay demonstration downlink at Table Mountain Facility[32].